### v3

# Synthetic Homes: An Accessible Multimodal Pipeline for Producing Residential Building Data with Generative AI

Abstract
Computational models have emerged as powerful tools for multi-scale energy modeling research at the building level as well as urban scale. However, these models require a plethora of data on building parameters, some of which can be inaccessible, expensive, or can raise privacy concerns. We introduce a modular multimodal framework to synthetically produce this data from publicly accessible images and residential information using generative Artificial Intelligence (AI). Additionally, we provide a modeling pipeline demonstrating this framework and we evaluate its generative AIcomponents for realism. Our experiments show that our framework’s use of AIavoids common issues with generative models and produces realistic multimodal data at the building scale. Resulting datasets can be used for assessing influence of energy efficiency upgrades at the building scale, as well as to simulate larger patterns of energy consumption across regions. This work will support research in building and energy simulation by reducing dependence on costly or restricted data sources, and pave a path towards more accessible research in Machine Learning (ML)and other data-driven disciplines.
Keywords: Building Energy Models (BEM), Urban Building Energy Modeling (UBEM), Energy Policy, Synthetic Data, Generative AI
1 Introduction
Consumption of electricity in the United States will increase by $1.7\%$ between 2020 and 2026 (U.S. Energy Information Administration, 2025) and seasonal electricity use per person has nearly doubled from 1973 to 2025 (U.S. Environmental Protection Agency, 2025). Rising energy consumption, due to inefficient building mechanics, exerts increasing pressure on infrastructure capacity, operational costs, and greenhouse gas emissions, resulting in rising costs for governments, energy producers, and energy consumers. Accurate modeling and policy tools are therefore essential for effective planning, aiding by preventing infrastructure strain, optimizing energy costs, minimizing or mitigating climate impacts, assessing policy effectiveness, and guiding decision-making in urban centers (Bompard et al., 2024; Perera et al., 2021; Zeng et al., 2011; Hajri et al., 2025).
Computational approaches—such as the use of Building Energy Models (BEMs)to monitor energy usage and model urban centers, buildings, and other structures—require large amounts of data, such as type and quality of materials, floor plans, and weather patterns to produce scalable and quantitative results (U.S. Department of Energy, ). However, several issues arise when attempting to collect such information, including high prices, limited availability, and privacy concerns. These issues create barriers to accessing data, hindering the development of accurate models and the scaling of models to larger populations or geographic areas. The inaccessibility of this data has led researchers to generate synthetic data through a variety of methods. We propose a generative AIbased approach to generate low-cost multimodal synthetic data consisting of simulations and home inspection notes that can be used to develop, validate, and improve existing computational models for energy demand modeling. The proposed method enables the generation of large volumes of data for use in any geographic region where basic descriptive data on housing structures is publicly available.
In the following sections, we position our work in the existing literature (Section 2) and provide an in depth view of our framework (Section 3). Then, we evaluate the realism of data generated by our pipeline (Section 4) using a dataset created to statistically model the region of our synthetic homes. Finally we conclude and highlight future directions from this work (Section 5).
2 Related Works
Current data generation efforts span a wide range of approaches, from optimizing building construction processes to the creation of digital twins. For example, Wan et al. (2022) built a generative model that improves the efficiency of energy consumption by determining the optimal layout of floor plans. Elnabawi and Hamza (2022) enhanced their building energy models using highly detailed microclimate models that considered features such as shadow generation and seasons to generate weather data. Lee et al. (2019) train a neural network that predicts building energy consumption, aiming to create a more robust model with limited real-world measurements. The network was trained with data synthesized using, among others, electricity, temperature, and humidity, encoded into receiver operating characteristic (ROC) plots which visually encode the separation between classes and are commonly used to assess model discrimination (Zweig and Campbell, 1993; Janssens and Martens, 2020).
Other researchers focus on creating and applying digital twins: virtual, dynamic, and data-driven digital replicas of real-world urban environments. Stinner et al. (2021) generate digital twins of building energy systems from piping and instrumentation diagrams using convolutional networks and other algorithms. Agostinelli et al. (2021) create digital twins of energy efficient buildings in residential areas from Building Information Modeling (BIM), Internet of Things (IoT), Geographical Information Systems (GIS), and AIcomponents, with machine learning serving as the central module of the model. Francisco et al. (2020) take a different approach by generating digital twins of daily, time-segmented energy benchmarks solely from smart meter data streams. Belik and Rubanenko (2023) present a hybrid digital twin framework for renewable energy sources (RES) that integrates real-time data, 3D modeling, and sensitivity analysis to improve forecasting and operational efficiency, particularly in grid systems with high RES penetration.
While prior work has focused on traditional simulation tools or sensor data to construct digital twins, Xu et al. (2024) encourage the use of generative AImodels, such members of the GPTfamily, because they present new opportunities in urban data generation. For instance, Dodge et al. (2017) employ convolutional neural networks (CNNs) to parse floor plan images, generating structured representations using wall segmentation, object recognition, and optical character recognition. Relatedly, Zhang et al. (2025a) implement a pipeline that generates EnergyPlusIDF files from initial building geometry and textual building descriptions. Moreover, Xiao and Xu (2024) propose a multi-agent Large Language Model (LLM)framework that processes unstructured building data and sensor logs to support automated energy optimization in smart grid contexts. Additionally, in closely related fields, generative AIhas been successfully employed to perform a variety of tasks, such as predictive analysis of traffic congestion (Lin et al., 2025), real-time decision-making in urban air mobility (Sha et al., 2024), and stochastic solar irradiance forecasting for building facades (Zhang et al., 2022).
Liu et al. (2025) present a wide-ranging review on the integration of generative AIinto building energy workflows, highlighting both opportunities and challenges. These include concerns such as dataset preparation, hallucinations, and a lack of domain-specific expertise. However, the impact of such challenges varies depending on the pipeline design. Because our framework combines pretrained LLMswith simulation tools for quantitative evaluation, we avoid the need for fine-tuning or physics-aware generation, thereby reducing the impact of these concerns. We evaluate our framework’s LLMcomponents to reduce hallucinations. Additionally, Liu emphasizes the importance of multimodal integration—where multiple data modalities (such as image, text, and tabular data) are used by a process or model. Producing high-quality structured urban data suitable for tasks such as modeling and simulation requires combining multiple complementary processes into a single workflow (Rehmann et al., 2025; Guo et al., 2023; Bishop et al., 2024). We adopt these principles in the design of our general framework.
3 Methodology
We propose a novel framework that combines multimodal real-world data with data generated using LLMs. We leverage open-access county data and generative AIto construct synthetic homes that mirror real-world counterparts and inspection notes corresponding to each. Home inspection notes will be used as a source to rate the homes based on their energy efficiency, alongside home images, floorplans, and simulation results. Finally, we test the LLMcomponents of our framework and the realism of the generated data. Our framework comprises four modular components, depicted in Figure 1, where each stage focuses on a transformation of the data, from data collection to simulation:
  1. 1.
    A web scraper that collects data and images;
  2. 2.
    An image processor powered by LLaVA (Liu et al., 2023);
  3. 3.
    A Geographic JSON (GeoJSON)and inspection note generator using GPT (OpenAI, 2025); and
  4. 4.
    An EnergyPlus (Crawley et al., 2001) simulation
Figure 1: Overview of the pipeline.
While our pipeline is tailored to generate data from images and tabular data, it is inherently modular and extensible. Any stage is easily modified or expanded for other related purposes, and different information can be generated in addition to or in place of any data produced by this pipeline. For example, a different LLMcould be employed in place of GPT, or other types of data could be considered or generated by the pipeline. Those interested in generating synthetic commercial buildings could use a modified version of this pipeline that emphasizes the characteristics and data central to their research. This opens use of our work to a larger set of individuals. Rather than proposing a rigid set of steps, we present a flexible framework that integrates generative AI, web scraping, and building simulation, designed for rapid, low-cost data generation in building energy and other fields that traditionally suffer from data scarcity. Ultimately, the data generated by our framework can be used to prototype, design, train, and research urban energy systems and models.
An instance of this pipeline was implemented and executed [^1] Our source code is available at: https://github.com/Lafayette-EshbaughSilveyra-Group/synthetic-homes. on a g3.medium flavor JetStream 2 virtual machine running Ubuntu 24.04.2 (Noble). Specifically, the system featured one NUMA node with 8 AMD EPYC-Milan processors, 30 GB of RAM, and an NVIDIA A100 SXM4 40 GB GPU (of which 10 GB was allocated). Additionally, it was fully virtualized (KVM and AMD-V). We paid \$0.36 (USD) for OpenAI requests to produce data for 258 homes—about \$0.0014 per home. [^2] Assuming a conservative upper bound of 10 CPU minutes per home at 100 watts and electricity prices at \$0.15 per kilowatt-hour, local electricity costs are approximately \$0.0025 per synthetic home, giving a worst case total of \$0.0039 per synthetic home. To run the pipeline, we used the PyTorch (Ansel et al., 2024) machine learning framework, and the Hugging Face Transformers library (Wolf et al., 2020) to load the models we used. In the subsections that follow, we describe each pipeline component in detail.
3.1 Scraping the Data
To ensure the synthetic data we generated reflects real-world homes, we collected data from publicly available county datasets. For our pipeline instance, we gathered data from [location withheld for double-blind review]. To collect this data, we implemented a web scraper using Selenium WebDriver (Selenium, 2024) and ChromeDriver (Google, 2023). For each residential property on each street in the list of streets to process, the scraper extracts key attributes, such as number of rooms, total floor area, and other metadata listed in Table 1. This data is saved to a JavaScript Object Notation (JSON)file for each home. Additionally, the scraper downloads two images: a street view photograph and a floor plan, both of which are made available on the county’s public-facing platform. An example of each is depicted in Figure 2.
Figure 2: An example street view photograph and floor plan from the county database.
Collected Data Type
Year Built Year
Remodeled Year Year
Land Use Code String
Total Square Feet Living Area Number
Number of Stories Number
Grade String
CDU String
Building Style String
Total Rooms Number
Bedrooms Number
Full Baths Number
Half Baths Number
Additional Fixtures Number
Total Fixtures Number
Heat/Air Cond String
Heating Fuel Type String
Heating System Type String
Attic Code String
Unfinished Area Number
Rec Room Area Number
Finished Basement Area Number
Fireplace Openings Number
Fireplace Stacks Number
Prefab Fireplaces Number
Basement Garage (Number of Cars) Number
Condo Level Number
Condo/Townhouse Type String
Basement String
Exterior Wall Material String
Physical Condition String
Sketch Data Dictionary of floorplan segment areas
Table 1: All data we collected from the county portal. Not all homes have values for each variable.
3.2 Processing Images
After scraping the street view photographs and floor plans from the county portal, they are analyzed using LLaVA (Liu et al., 2023). This step processes both files to translate them into a textual description. For example, the floor plans can be used to define the geometry of the home, and the image to determine the quality and number of windows. The use of images allows our pipeline to infer important aspects of the homes that will be used to generate the GeoJSON, which is ultimately used to generate the Input Definition File (IDF)file required by EnergyPlus.
3.2.1 Evaluating Visual Focus with Occlusion Testing
In image processing, it is crucial that only key portions of the images in question trigger differences in the response of a model. We refer to this as focus. For example, if the prompt asks about the state of the roof of a home, the tree next to the home should be inconsequential in comparison to the roof itself. To evaluate this, we use a technique called occlusion.
Occlusionmeasures the necessity of a part of an image: if removing a feature causes the prediction to change, that feature was necessary for the model’s decision (Hooker et al., 2019; Balkir et al., 2022). When we remove image portions, we replace them with a neutral value, such as black. This creates a masked version of the image. The degree of change measured between the model’s output for this masked image and its baseline response reflects the importance of the masked cell. High sensitivity to the removal of a feature suggests the model relied heavily on that region to produce its output, indicating that the region was internally necessary to the model’s decision process.
During the development of our pipeline, we identified two candidate LLMsto employ for image processing. To determine the best model for this task, we compared the focus of the GPTand LLaVALLMswhen analyzing images using occlusion tests. We provided the models with twenty different images of homes: ten with roofs showing no signs of damage, and ten with roofs with significant damage. Each model was instructed to evaluate and describe the state of the roof and justify its evaluation. We first took a baseline where no part of the image was masked. Then, the image was divided into one hundred cells, a number we selected to balance effectiveness, efficiency, and cost. Each cell of the image was masked, and the model was executed on the original prompt and this new image. Textual responses were embedded into vectors using the all-MiniLM-L6-v2 SentenceTransformers model (Reimers and Gurevych, 2019), a lightweight and widely used model sufficient for obtaining sentence-level numerical representations. Then, to measure the semantic difference between each execution of the experiment, we took the cosine distance of each embedded response from the baseline. Additionally, heatmaps of these differences were assembled using matplotlib (The Matplotlib Development Team, 2025).
The mean difference value in roof cells and non-roof cells for each image was calculated. Figure 3 depicts the distribution of the forward occlusion results. These scatter plots are particularly telling—GPT(Figure 3(a)) evenly scatters the data points on either side of the line $y=x$, indicating it has no preference for which area in the image is considered important. In comparison, LLaVA(Figure 3(b)) tends to produce results in the upper half of the graph. This illustrates LLaVA’s improvement over GPTplainly: more often than not, LLaVA’s response depends more on the roof than other parts of the image.
(a) GPTresults.
(b) LLaVAresults.
Figure 3: GPTand LLaVAforward occlusion per image results.
Additionally, overall means and standard deviations were calculated for GPTand LLaVA(Figure 4). In a well-focused model, the mean difference in the relevant area should be greater than the mean difference anywhere else. This behavior occurs since relevant cells should account for the majority of the response given by the model. In our experiments, we observe that GPTreports identical roof and non-roof mean differences (Figure 4(a)). However, LLaVAachieved a roof mean difference approximately $20\%$ greater than its non-roof mean difference (Figure 4(b)).
Mean Difference SD
Roof 0.0956 0.0384
Non-Roof 0.0956 0.0408
(a) GPTocclusion overall statistics
Mean Difference SD
Roof 0.0208 0.0241
Non-Roof 0.0149 0.0197
(b) LLaVAocclusion overall statistics
Figure 4: Occlusion statistics for GPTand LLaVA.
When we compared individual GPTand LLaVAocclusion results, we found more evidence that LLaVAperforms better at focusing on key parts of the images. LLaVAis much more particular about which cells are key to construct the response; GPT produces results that look stochastic. Figure 5 displays one noteworthy example—LLaVAfocuses on a more specific section of the image, primarily consisting of the home’s roof. GPT, on the other hand, behaves much more randomly. Each value in the cell is the semantic difference, and the background tint of the cell represents the scaled semantic difference, where the brightest background represents the largest difference and the darkest represents the smallest. In Figure 5(a), the semantic difference is generally larger on the roof than on other parts of the image. A small semantic difference is a signal that a given portion of the image is less important to the model in order to formulate a response. In Figure 5(b), the majority of cells are moderately important to the model in formulating a response—even cells containing walls or grass. Comparing the two, LLaVAdemonstrates better focus on the more important parts of the image. Full occlusion experimental details, hypotheses, and findings are available in A.
(a) LLaVAocclusion results. These results show less sensitivity and more focus on the more important parts of the image.
(b) GPTocclusion results. These results demonstrate extreme sensitivity to the image, which is not desired. This sensitivity means the model focuses less on the most important parts of the image while responding to the prompt.
Figure 5: GPTand LLaVAocclusion results.
3.3 Generating GeoJSON
Once the images are processed, we utilize the descriptions given by LLaVAin a GPTprompt (see Appendix B) to generate a GeoJSONfile and home inspection notes. The prompt provides the image descriptions (Section 3.2) and data that was scraped from the county portal (Section 3.1) and directs the model to take two actions:
  • The prompt instructs GPTto generate a GeoJSONfile for the building, including geometry, data from the county, and five estimated performance parameters. These estimated parameters, defined according to standard convention, include the HVAC heating and cooling coefficients of performance, r-values for the roof and wall, and the air change rate.
  • The prompt directs GPTto write a short inspection note, focused on energy-related observations such as insulation, HVAC type/age, visible windows, and any inferred upgrades. This inspection note is placed in the GeoJSONfile.
The prompt contains strict guidelines to ensure outputs are valid, accurate, and complete. An excerpt from an example GeoJSONobject is provided in Figure 6.
We use GPTfor structured data generation given its strong ability to produce well-formed JSONoutputs from descriptive prompts. Multiple studies have shown that LLMsperform well on generation tasks under strict schema constraints, even with no previous prompting or fine tuning. For example, StructuredRAG’s benchmark suite showed $82\%$ format compliance under JSONprompt conditions (Shorten et al., 2024). As we built the pipeline, we reviewed GPT’s responses and it seldom failed to adhere to the JSONstructure. In the rare case that it doesn’t, our program catches the JSONerror and automatically reprompts GPTfor that example.
1{
2 "type": "Feature",
3 "properties": {
4 "name": "Generated Home",
5 "floor_area": 2576,
6 "building_type": "Single family",
7 "inspection_note": "...",
8 "hvac_heating_cop": 0.85,
9 "hvac_cooling_cop": 3.0,
10 "wall_r_value": 13,
11 "roof_r_value": 30,
12 "air_change_rate": 0.35
13 },
14 "geometry": {
15 "type": "Polygon",
16 "coordinates": [...]
17 }
18}
Figure 6: Excerpt of a generated GeoJSONfile.
3.4 Running EnergyPlus Simulations
To execute a simulation in EnergyPlus, we take the GeoJSONgenerated by GPT(Section 3.3) and convert it to an IDFfile for use in EnergyPlus. We use a template IDFfile, filled with default values that are replaced with variables from the GeoJSONdata, such as the geometry, HVAC heating and cooling coefficients of performance, r-values for the wall and roof, and the air change rate. Additionally, the template has many fields with constant values, such as the simulation running period and materials, constructions, and zones used in the simulation. Once the details from the GeoJSONare copied into the template, a final IDFfile is produced.
The templating process is facilitated with the eppy package (santoshphilip, 2024), which makes programmatic IDFgeneration straightforward. Then, we use ExpandObjects to prepare the file for simulation. ExpandObjects is bundled with EnergyPlus, and its primary use case is to take an IDFand expand template objects—such as the HVACTemplate we use—into full systems (U.S. Department of Energy, 2023). Finally, we run the expanded IDFfile through EnergyPlus, producing final simulation results.
We write the inspection notes and EnergyPlus simulation results for each example into one JSON Lines (JSONL)file. This completes our end-to-end dataset generation process, yielding a unified dataset that combines textual and simulation-based information.
4 Validating Realism of Pipeline Generated Data
Following prior work on synthetic data evaluation (Goncalves et al., 2020; Patki et al., 2016), we adopt the definition of realism as the overlap of synthetic values with the $10\%$$90\%$ range of a trusted reference on the home characteristics being generated. To measure the realism of our generated data, we compare it with ResStock (Present et al., 2024; Reyna et al., 2025), a dataset that contains realistic samples of homes in different regions of the United States as IDFfiles. The analysis consists of performing statistical analyses comparing the distribution of values in our synthetic homes to those in ResStock. We only use ResStock data from climate zone 4A, the region in which the synthetic homes are collected in our example, and we only compare the four pre-selected variables—Wall R-Value, Roof R-Value, Cooling Coefficient of Performance (COP), Heating COP—that are not static among all synthetic homes. The number of synthetic homes is $258$ while the number of climate zone 4A ResStock samples is $2000$.
Table 2 reports the realism for the four variables across all synthetic homes by displaying the percentage of values that overlap with the $10\%$$90\%$ range of ResStock. We found overlap greater than $65\%$ for all variables and greater than $90\%$ for three variables. This overlap is a strong indicator of realism for our generated data. Additionally, the table includes the mean and median of the four variables for all synthetic homes and the interquartile range (IQR) of the four variables for all ResStock homes. These statistics are included for informational purposes and are not considered when determining realism.
Variable Overlap (%) ResStock IQR (Q1–Q3) Synthetic Mean Synthetic Median
Wall R-Value 93.41 0.704 (1.94–2.64) 2.79 2.29
Roof R-Value 100.0 3.35 (3.35–6.69) 5.04 5.28
Cooling COP 66.28 0.70 (3.1–3.8) 2.99 3.00
Heating COP 100.0 0.20 (0.8–1.0) 0.84 0.85
Table 2: Coverage of synthetic homes within the 10–90% ResStock range, alongside ResStock interquartile ranges (Q1–Q3) and synthetic homes mean and median.
The results above and the tests of the generative components of our framework demonstrate a level of confidence above previous work such as Zhang et al. (2025a) and Xiao and Xu (2024). These works and others either accept the output of the LLMsor use procedural evaluation techniques.
Occlusion offers stronger empirical support that the image processing component of our pipeline—which is LLM-based—performs as expected. We additionally demonstrated statistically that properties of our generated homes are realistically distributed according to ResStock. Our results directly address and mitigate concerns about hallucination and unreliability raised in prior work (Liu et al., 2025). It is fair to be skeptical of LLMsand other MLmodels—but with our results, we clearly demonstrate our pipeline is not as susceptible to some of the issues Liu et al. (2025) raise.
5 Conclusion
In this work, we presented a novel way to generate multimodal building-scale data for urban energy research and other related fields. Using a multimodal pipeline, we produced this data in a cost-effective and efficient way. We also performed an in-depth analysis to ensure that our generative AIpipeline components were behaving as expected, taking into account the concerns of Liu et al. (2025).
Generative AIis a powerful tool, and we recognize the importance of this new frontier as described by Xu et al. (2024) and  Liu et al. (2025). Not only does our work demonstrate a step forward in the urban energy field, it also presents a useful, malleable tool that experts and researchers can use for a variety of tasks.
One promising direction is the use of this pipeline to train MLalgorithms. A natural expansion of the pipeline could utilize the resulting EnergyPlus simulation data and inspection notes to rank the efficiency of particular components that affect the energy efficiency of a home. These labeled homes could be used as inputs for an MLmodel that could provide recommendations for energy efficiency retrofits. Furthermore, this model, trained with our synthetic data, could be used to predict the priorities for retrofits in real houses.
Overall, the high prices, lack of availability, or privacy concerns attached to the collection of urban energy data can be avoided by generating synthetic data. With our framework, all of these concerns are alleviated and realistic data is produced. This reduced dependence on costly or restricted data sources provides a pathway towards more accessible research.
Appendix A Occlusion Results
In this appendix, we provide detailed information and results from our occlusion tests. Firstly, when running the test we utilized identical prompts for each model. For GPT, we prompted:
You are a certified home inspector. Describe the status roof. Is it in good condition? Why or why not?
Likewise, we prompted LLaVA with the following:
USER: <image>
You are a certified home inspector. Describe the status roof. Is it in good condition? Why or why not? ASSISTANT:
These prompts differ slightly due to how the models are trained; namely, LLaVA expects an image token and labeled portions of the message. The majority of LLMs behave this way, and the GPT API likely wraps the given prompt in something similar.
Next, we provide our hypotheses pertaining to the performance of both models. We hypothesize that LLaVA’s advantage stems from its architecture: it employs an early fusion design in which images are converted into visual tokens by CLIP (Radford et al., 2021) that are then prepended to the textual prompt (Liu et al., 2023). These visual tokens are passed through the transformer layers of the LLMtogether with text tokens, enabling multimodal attention at every layer. It has been demonstrated that vision tokens are most critical for a model in early layers, when they fuse into text tokens (Zhang et al., 2025b). A likely result of this is the reduction of the model’s sensitivity to variations in the image, which may explain its stronger performance in our experiments.
While GPT-4V (GPT’s vision model) is proprietary, the case can be made that it possesses a different structure. Evaluation studies show that it performs well across visual reasoning and captioning tasks, but it falls short in fine grained visual problems, requiring complex visual tasks to be broken up into interleaved subfigures and text (Yang et al., 2023). These evaluations are similar to evaluations from late fusion (connector) models. For example, BLIP-2 shows strengths in captioning images and instruction-following, while it is less robust with fine-grained spatial reasoning (Li et al., 2023). Another model, MiniGPT-4, demonstrates a similar profile: it has strengths in tasks such as image description, but falters with complex spatial reasoning (Zhu et al., 2023). Both BLIP-2 and MiniGPT-4 are late fusion models, where compressed visual features are routed into the LLMat later stages (Li et al., 2023; Zhu et al., 2023). As shown, BLIP-2, MiniGPT-4, and GPT-4V all suffer from a very similar weakness: fine-grained or complex spatial reasoning. Based on this similarity, we hypothesize that GPT-4V is also a late fusion model. This helps explain why GPTperformed differently from LLaVAin our experiments: as structure dictates function, differences in behavior strongly suggest differences in design. On the basis of this analysis, we feel confident to select LLaVAto complete this task.
In Table 3, we provide our forward occlusion results, including roof mean difference (RMD) and non-roof mean difference (NRMD) values from both GPT and LLaVA and references to figures displaying GPT and LLaVA heatmaps side by side.
Metric 1 2 3 4 5 6 7 8 9 10
Damaged Roofs
GPT RMD 0.0856 0.0920 0.0730 0.0689 0.1127 0.0453 0.0516 0.0603 0.1149 0.0419
GPT NRMD 0.0903 0.0745 0.0615 0.0824 0.0900 0.0375 0.0485 0.0900 0.1000 0.0442
LLaVA RMD 0.0033 0.0324 0.0949 0.0093 0.0262 0.0269 0.0356 0.0055 0.0038 0.0208
LLaVA NRMD 0.0057 0.0545 0.0723 0.0059 0.0200 0.0000 0.0335 0.0055 0.0022 0.0042
Roof Cells 70 80 53 83 97 96 80 96 91 88
Non-Roof Cells 30 20 47 17 3 4 20 4 9 12
Figure 7 8 9 10 11 12 13 14 15 16
Undamaged Roofs
GPT RMD 0.1128 0.0785 0.1465 0.1508 0.0822 0.0956 0.1087 0.2005 0.0696 0.1200
GPT NRMD 0.1055 0.0742 0.1872 0.1598 0.0884 0.0930 0.1230 0.1898 0.0768 0.0959
LLaVA RMD 0.0709 0.0085 0.0007 0.0000 0.0340 0.0108 0.0066 0.0000 0.0017 0.0240
LLaVA NRMD 0.0313 0.0062 0.0000 0.0000 0.0321 0.0023 0.0030 0.0000 0.0015 0.0181
Roof Cells 53 47 68 49 81 39 70 60 47 73
Non-Roof Cells 47 53 32 51 19 61 30 40 53 27
Figure 17 18 19 20 21 22 23 24 25 26
Table 3: Quantitative results from our forward occlusion experiments, including roof mean difference (RMD) and non-roof mean difference (NRMD) for both GPT and LLaVA. The number of cells in the image considered roof and non-roof are listed, and figures of side by side comparisons of the resulting heatmaps are referenced.
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 7: Forward Occlusion—Damaged Roof 1
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 8: Forward Occlusion—Damaged Roof 2
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 9: Forward Occlusion—Damaged Roof 3
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 10: Forward Occlusion—Damaged Roof 4
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 11: Forward Occlusion—Damaged Roof 5
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 12: Forward Occlusion—Damaged Roof 6
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 13: Forward Occlusion—Damaged Roof 7
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 14: Forward Occlusion—Damaged Roof 8
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 15: Forward Occlusion—Damaged Roof 9
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 16: Forward Occlusion—Damaged Roof 10
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 17: Forward Occlusion—Undamaged Roof 1
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 18: Forward Occlusion—Undamaged Roof 2
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 19: Forward Occlusion—Undamaged Roof 3
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 20: Forward Occlusion—Undamaged Roof 4
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 21: Forward Occlusion—Undamaged Roof 5
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 22: Forward Occlusion—Undamaged Roof 6
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 23: Forward Occlusion—Undamaged Roof 7
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 24: Forward Occlusion—Undamaged Roof 8
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 25: Forward Occlusion—Undamaged Roof 9
(a) GPT Heatmap
(b) LLaVA Heatmap
Figure 26: Forward Occlusion—Undamaged Roof 10
Appendix B GeoJSON Generation Prompt
In this appendix, we provide the prompt used to generate the GeoJSON data in Section 3.3.
You are a certified home energy inspection expert and data specialist
building synthetic training data for an AI model.
You are provided with:
- Structured residential property data (JSON).
- A detailed exterior description of the home: "{exterior_description}"
- A detailed floorplan description: "{floorplan_description}"
Your tasks:
1. Generate a GeoJSON file for this building with:
- A plausible (longitude, latitude) location in Bethlehem, PA.
- A FeatureCollection containing exactly one Feature.
- Geometry: Polygon or MultiPolygon representing the building footprint,
realistic for the reported square footage and building style.
- Properties from the provided JSON:
- Year Built
- Total Square Feet Living Area
- Building Style
- Exterior Wall Material
- Heating Fuel Type
- Heating System Type
- Heat/Air Cond
- Bedrooms
- Full Baths
- Half Baths
- Basement
- Number of Stories
- Grade
- Estimated performance parameters:
- hvac_heating_cop (0--1)
- hvac_cooling_cop
- wall_r_value
- roof_r_value
- air_change_rate (0--1)
2. Write a short inspection note as if you had toured the home,
focusing on energy-related observations.
Strict Guidelines:
- Only base outputs on the provided data and descriptions.
- Do not invent unsupported details.
- Ensure the GeoJSON is valid and realistic.
- Coordinates must plausibly place the home in Bethlehem, PA.
Structured property data:
{json.dumps(house_data)}
Output format:
"geojson": {
"type": "FeatureCollection",
"features": [
"type": "Feature",
"geometry": { ... },
"properties": {
"air_change_rate": ...,
"hvac_heating_cop": ...,
"hvac_cooling_cop": ...,
"wall_r_value": ...,
"roof_r_value": ...
"inspection_note": "..."
No backticks or explanation.
Glossary
\@glotype@acronym@title